# CALRec: Contrastive Alignment of Generative LLMs For Sequential Recommendation

### v1

@@ 6. Further Analyses and Discussion
W/O Dedup - Metric (opt) W/ Dedup - Metric (opt)
@@
Table 6. Data deduplication. Optimistic evaluations of LIR Baseline (always recommend last-item in the input sequence) on data from (Li et al., 2023b) (w/o dedup) and our data w/ dedup respectively.
Output Matches an Item (%) Hierarchical Match - Metric: (opt)/(pes) & Improv. vs BM25)
@@
Table 7. Hierarchical match results. The relative gains/drops on $\frac{\text{opt}+\text{pes}}{2}$ are also reported against BM25 retrieval.
Further Discussion on User Sequence Dedup. Now, we further delve deeper into user sequence dedup issues. Although in real world users can make repeated activities, we still posit that the high ratio of consecutive identical interactions in the raw Amazon Review data are not likely to reflect normal/real users’ behaviors (further details in Appendix). While we cannot conclude that data deduplication is ‘correct’ or ‘wrong’, we reckon that the ‘LIR’ baseline results are meaningful and should be reported for model comparison purposes. We further run the LIR Baseline on 1) data from Li et al. (2023b) without dedup and compare with its results on 2) our data with dedup. Although the data statistics are slightly different, the comparisons are still meaningful, since data sizes are both considerable and thus are representative of the full distribution. Table 6 shows the optimistic metrics of LIR Baseline that always recommends last item in the input sequence on data w/o dedup (left four columns) and data w/ dedup (right four columns). All metrics suffer sharp reductions due to the dedup process. Specifically, Recall@1 drops $60\%$-$95\%$ after removing consecutive identical events. We emphasize the necessity of sequence interaction deduplication as it is critical to reliably measuring model quality.
@@
First, we found that almost all ($>99.9\%$) finetuned model output honors the desired item text format as “Title: **. Category: **. Brand: **. Price: **.”. We then extract the four attribute entities: title, brand, category and price, and find that $82\%-98\%$ of model outputs can find an item in the item corpus that matches in all four attributes, and $95\%-99.8\%$ of title entities extracted from the model outputs exists in the item corpus. These statistics are summarized in Table 7. Based on these findings, we propose a heuristic hierarchical matching method. We first search exact matches in the item corpus for all four attributes extracted from model outputs; If no match is found, we search for the exact match of the first three attributes, then the first two, and finally only the title description. Similar to our main experiments, both optimistic and pessimistic metrics are calculated and results are also shown in Table 7: we find that our hierarchical matching method has marginally worse performance compared to the BM25-based retrieval but still outperforms other baselines models shown in Table 2.
Limitations and Future Work. Due to the fact that fine-tuned CALRec usually outputs exact item descriptions seen during training, a direct limitation ensues: CALRec barely cannot address cold-start scenarios where the ground-truth item is not seen in the training nor in the validation set. We leave addressing this limitation to future work.
Second, although CALRec outperforms other baselines in Recall@1, NDCG@10, and MRR@$\infty$ by large margin, its Recall@10 underperforms one of the text+ID combined baseline. We attribute this to the fact that temperature sampling, unlike embedding-based nearest neighbor search, is a quasi-greedy ranking method and may not yield the best top 10 predictions among the entire item corpus. Therefore, although CALRec performs well in metrics that emphasize the order of top predictions (like Recall@1 and MRR), its top-10 retrieval prediction needs more diversity.
Third, in real-world applications, when the size of item corpus is extremely large (e.g., billions of items), BM25 retrieval becomes extremely slow and even infeasible. In addition to the hierachical entity matching method discussed in the previous paragraph, we may extend CALRec to a retrieve-and-re-ranking model where any baseline sequential recommendation approach can be used to first derive, for example, top 10K recommendations focused on high recall. We can then make use of CALRec’s output texts to rerank these items which can not only leverage the strong capabilities of CALRec but also avoid doing BM25 retrieval on an extremely large corpus.
@@ 7. Conclusion
We proposed CALRec, a contrastive-learning-assisted two-stage training framework for sequential recommendation based on LLMs, where we run experiments with PaLM-2 LLM backbone. Our approach also features careful template design inspired from few-shot learning and a novel quasi-round-robin BM25 retrieval approach. Through comprehensive experiments on the Amazon Review dataset, we demonstrated the effectiveness of the proposed method supported by PaLM-2, where we outperform existing SotA models for sequential recommendation by a considerable margin in most evaluation metrics. We also presented a series of further analyses and revisited the user sequence duplication issue and existing evaluation metrics; we accordingly propose to adopt a training-free text-statistic based ‘Last Item Repeater’ reference baseline for model comparison purposes, and we strongly suggest reporting both optimistic and pessimistic metric scores for all text-based approaches.
@@ Appendix A Examples of Data Duplication
As discussed in §4, the raw amazon review datasets contain repeated, non-distinguishable user-item interactions that share identical entity values across all attributes, including item id, user id, timestamp, item metadata, user rating, user review text, etc. An example in the industrial and scientific category is as follows:
@@
In the example provided, all purchase-review events, except the initial one, are identical. This behavior generally deviates from normal user activities and may potentially confuse or mislead models into recommending repetitive items to achieve artificially inflated metrics. Although only a small percentage of the events are duplicates (Table 1), our analysis via the LIR baseline on raw and dedup datasets (Table 6) shows that those duplicates can have a significant impact on the behavior of the models because they provide models with an easy, oversimplified prediction path.
Appendix B Effectiveness of Contrastive alignment
Fig 2 shows the loss profiles during category-specific finetuning on the Scientific category without first multi-category joint fine-tuning. The left panel shows the model start overfitting in text generation task at $100$ epochs in the industrial and scientific category. The right panel shows that train, evaluation and test contrastive loss between prediction and target drops quickly from $6.2$ to $4.4$ and remains relatively stable. At batch size $512$, the initial $\mathcal{L}_{TT}\approx 6.2$ means the average ratio between $\exp(\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{i}^{T})/\tau_{c})$ and $\exp(\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{j}^{T})/\tau_{c}),j\neq i$ is about $1.04512\times\exp(-6.2)=$ 1.04 at the beginning of the training. Thus $\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{i}^{T})-\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{j}^{T})=\ln(1.04)\tau_{c}\approx 0.02\ll 1$, suggesting that without finetuning, the raw pretrained LLMs, although able to make some prediction (Tab 5), does not yield prediction embeddings that are more aligned with the ground truth items than random items. After finetuned and $\mathcal{L}_{TT}\approx 4.4$, the same ratio increases to $6.3512\times\exp(-4.4)=$ 6.3, corresponding to $\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{i}^{T})-\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{j}^{T})\approx 0.92\sim 1$, indicating that the predicted item embedding is $\sim 90^{\circ}$ more aligned to target embedding than a random item embedding. On the contrary, without contrastive alignment, the prediction-target contrastive loss $\mathcal{L}_{TT}$ only drops from $6.4$ to $6.0$ as the model learns to generate the correct next-item prediction as a pure text generation task. Fig 2 demonstrates that the contrastive loss can efficiently align the model prediction embedding and the target embedding.
Figure 2. Text generation cross-entropy loss ($\mathcal{L}_{NIG}$, left panel) and prediction-target contrastive loss ($\mathcal{L}_{TT}$, right panel) during Scientific-category-specific finetuning without multi-category joint fine-tuning. In both panels, the training curves (red-dotted) without contrastive alignment are also plotted for comparison.

### v2

@@ 6. Further Analyses and Discussion
w/o Dedup - Metric: (opt) w/ Dedup - Metric: (opt)
@@
Table 7. Data deduplication. Optimistic evaluations of LIR Baseline (always recommend last-item in the input sequence) on data from (Li et al., 2023c) (w/o dedup) and our data w/ dedup respectively.
Output Matches an Item (%) Hierarchical Match - Metric: (opt)/(pes) & Improv. vs BM25 (%)
@@
Table 8. Hierarchical match results. The relative gains/drops in $\frac{\text{({opt})}+\text{({pes})}}{2}$ are also reported against BM25 retrieval.
Further Discussion on User Sequence Dedup. Now, we further delve deeper into the issue of duplicate interactions in user sequences. Although in real world users can make repeated activities, we still posit that the high ratio of consecutive identical interactions in the raw Amazon Review data are not likely to reflect normal/real user behavior (further details in Appendix/Auxiliary Materials). While we cannot conclude that data deduplication is ‘correct’ or ‘wrong’, we reckon that the ‘LIR’ baseline results are meaningful and should be reported for model comparison purposes. We further run the LIR Baseline on 1) data from Li et al. (2023c) without dedup and compare with its results on 2) our data with dedup. Although the data statistics are slightly different, the comparisons are still meaningful, since their data sizes are comparable and they are both representative of the full distribution. Table 7 shows the optimistic metrics of LIR Baseline that always recommends last item in the input sequence on data w/o dedup (left four columns) and data w/ dedup (right four columns). All metrics suffer sharp declines due to the dedup process. Specifically, Recall@1 drops $60\%$-$95\%$ after removing consecutive identical events. This highlights the importance of user sequence interaction deduplication as a critical step for reliably measuring model quality.
@@
First, we found that nearly all ($>99.9\%$) fine-tuned model outputs adhere to the desired item text format: “Title: **. Category: **. Brand: **. Price: **.”. We then extract the four attribute entities: title, category, brand, and price, and find that $82\%-98\%$ of model outputs correspond to an item in the item corpus that exactly matches all four attributes, and $95\%-99.8\%$ of title entities extracted from the model outputs exist in the item corpus. These statistics are summarized in Table 8. Based on these findings, we propose a heuristic hierarchical matching method. We first search exact matches in the item corpus for all four attributes extracted from model outputs; If no match is found, we search for the exact match of the first three attributes, then the first two, and finally only the title description. Similar to our main experiments, both optimistic and pessimistic metrics are calculated, with results also presented in Table 8: we find that our hierarchical matching method has marginally worse performance compared to our BM25-based retrieval, yet it still outperforms other baselines models presented in Table 2.
Limitations and Future Work. Due to the fact that fine-tuned CALRec usually outputs exact item descriptions seen during training, a direct limitation ensues: CALRec barely cannot address the item cold-start scenario where the ground-truth item is neither covered in the training set nor in the validation set. We leave addressing this limitation to future work. While previous approaches that also rely on matching LLM-generated item descriptions (e.g., titles) did not provide experimental results on item cold-start cases (Ji et al., 2024; Li et al., 2023d), we emphasize that our work is the first to identify this issue, which may also be present in these and similar approaches (yet to be verified). We regard this finding a minor contribution of our study.
Second, although CALRec outperforms other baselines in Recall@1, NDCG@10, and MRR@$\infty$ by large margin, its Recall@10 underperforms one of our ID+text baselines. We attribute this to the quasi-greedy nature of temperature sampling (especially when adopting a relatively small temperature value). Unlike embedding-based nearest neighbor search, our approach may not consistently produce the most optimal top-10 item predictions from the entire item corpus. Consequently, although CALRec excels in metrics that emphasize the relevance of the very top item prediction (like Recall@1 and MRR), there is a need for greater diversity in its top-10 item predictions.
Third, in real-world scenarios, when the size of item corpus is extremely large (e.g., billions of items), BM25 retrieval becomes prohibitively slow and often infeasible. In addition to the hierachical entity matching method discussed earlier in this section, we may leverage CALRec as a reranker within a retrieve-and-rerank framework where any baseline sequential recommendation approach can be used to first derive, for example, top 10K recommendations focused on high recall. Then CALRec’s output texts are used to rerank these items. This retrieve-and-rerank paradigm can not only leverage the strengths of CALRec but also circumvents the need for BM25 retrieval on a massive corpus.
@@ 7. Conclusion
We proposed CALRec, a contrastive-learning-assisted two-stage training framework for sequential recommendation based on LLMs, with experiments conducted using the PaLM-2 LLM as the backbone. Our approach also features careful template design inspired from few-shot learning and a novel quasi-round-robin BM25 retrieval approach. Comprehensive experiments on the Amazon Review dataset demonstrate the effectiveness of CALRec, supported by PaLM-2, where it significantly outperforms existing SotA models for sequential recommendation by a considerable margin in most evaluation metrics. We also presented a series of further analyses and revisited the issue of duplicate interactions in user sequences and existing evaluation metrics; we accordingly propose to adopt a training-free text-statistic based ‘Last Item Repeater’ reference baseline for model comparison purposes, and we strongly suggest reporting both optimistic and pessimistic metric scores for all text-based approaches.
Acknowledgements.
We thank the following Google colleagues (in alphabetical order) for their technical advice and helpful discussions on this work: Aahil Mehta, Ao Liu, Boaz Cogan, Jianmo Ni, Minh Pham, Nicolas Aagnes, Wang-Cheng Kang, Xin Zhang, Yan Zhu. We also extend our gratitude to anonymous reviewers for their constructive feedback.
@@ Appendix A Examples of Data Duplication
As discussed in §4, the raw Amazon Review data contain repeated, non-distinguishable user-item interactions that share identical entity values across all attributes, including item id, user id, timestamp, item metadata, user rating, user review text, etc. An example in the ‘Scientific’ category is as follows:
@@
In the example provided, all $5$ user-item interactions, except the initial one, are identical. This behavior generally deviates from normal user activities and may potentially confuse or mislead models into recommending repetitive items. Although duplicate events make up only a small percentage of the dataset (Table 1), our analysis via the LIR baseline on raw and deduplicated data (Table 7) shows that those duplicates can significantly affect model behavior by offering an overly simplistic pattern to learn from.
Figure 2. Text generation cross-entropy loss ($\mathcal{L}_{NIG}$, left panel) and prediction-target contrastive loss ($\mathcal{L}_{TT}$, right panel) during Scientific-category-specific fine-tuning without multi-category joint fine-tuning. In both panels, the training curves (red-dotted) without contrastive alignment are also plotted for comparison.
Appendix B Additional Discussion on Contrastive Alignment
Fig. 2 shows the loss profiles during category-specific fine-tuning on the Scientific category without Stage I multi-category joint fine-tuning. The left panel shows that overfitting happens in text generation task after circa $100$ epochs. The right panel shows that train, evaluation and test contrastive loss between prediction and target drops quickly from $6.2$ to $4.4$ and remains relatively stable. At batch size $512$, the initial $\mathcal{L}_{TT}\approx 6.2$ means the average ratio between $\exp(\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{i}^{T})/\tau_{c})$ and $\exp(\cos(\mathbf{v}_{j}^{T|U},\mathbf{v}_{i}^{T})/\tau_{c}),i\neq j$ is about $1.04512\times\exp(-6.2)=$ 1.04 at the beginning of the training process. Thus $\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{i}^{T})-\cos(\mathbf{v}_{j}^{T|U},\mathbf{v}_{i}^{T})=\ln(1.04)\tau_{c}\approx 0.02\ll 1$, suggesting that without fine-tuning, the raw pretrained LLMs, although able to make some prediction (Table 5), does not yield embeddings for predicted items (in the user tower) that are more aligned with embeddings of the ground truth items (in the item tower) than the embeddings from random users do. After fine-tuning, $\mathcal{L}_{TT}\approx 4.4$ and the same ratio increases to $6.3512\times\exp(-4.4)=$ 6.3, corresponding to $\cos(\mathbf{v}_{i}^{T|U},\mathbf{v}_{i}^{T})-\cos(\mathbf{v}_{j}^{T|U},\mathbf{v}_{i}^{T})\approx 0.92$ which is close to $1$, indicating that the embeddingsare aligned very well geometrically. On the contrary, without contrastive alignment, the prediction-target contrastive loss $\mathcal{L}_{TT}$ only drops roughly from $6.4$ to $6.0$ as the model learns to generate the correct next-item prediction as a pure text generation task. Fig. 2 demonstrates that the contrastive loss can efficiently align the embeddings from our two towers.